OpenStudy (anonymous):

Which of the following is a rational number? square root 97square root 98, square root 99, square root 100 square root 97 square root 98 square root 99 square root 100

Elsa213 (elsa213):

What do you think the answer is?

OpenStudy (anonymous):

i think the answer is 98 @Elsa213

OpenStudy (acxbox22):

http://www.mathsisfun.com/rational-numbers.html

OpenStudy (cwrw238):

sqrt 98 - sqrt2 8 sqrt 49 = 7 sqrt 2 which is irrational

OpenStudy (cwrw238):

* that 8 should be *

OpenStudy (anonymous):

umm i think its 98 or 100 @cwrw238

OpenStudy (anonymous):

@acxbox22 am i correct with the answer 98?

OpenStudy (anonymous):

bc i believe 98 can be a fractional number

Elsa213 (elsa213):

I think u r incorrect.

OpenStudy (anonymous):

i think its that or 100

OpenStudy (cwrw238):

no its not sqrt 98 because sqrt2 is irrational

OpenStudy (anonymous):

oh ok

OpenStudy (acxbox22):

what is the square root of 100?

OpenStudy (cwrw238):

we are not sealing with 98 but sqrt98

OpenStudy (cwrw238):

* dealing

OpenStudy (anonymous):

actully i think its 97

OpenStudy (acxbox22):

dont guess what is the square root of 100?

OpenStudy (anonymous):

100

OpenStudy (cwrw238):

you are confused about these numbers these are all square roots we are looking at sqrt 97 NOT 97 itself

OpenStudy (cwrw238):

square root of 100 = 10 10 * 10 = 100 10 is rational

OpenStudy (acxbox22):

as in\[\sqrt{97}~and~\sqrt{98}~and~ \sqrt{99}~and~\sqrt{100}\]

OpenStudy (anonymous):

ok i dont get this like u idk what sqrt is or any of that

OpenStudy (acxbox22):

the answer is the sqrt of 100

OpenStudy (cwrw238):

square root of a number is another number which when multiplied by itself gives the original number examples square root of 16 is 4 because 4 times 4 = 16 square root of 81 is 9 because 9 * 9 = 81

OpenStudy (anonymous):

oh ok i get it thats pretty simpal

OpenStudy (cwrw238):

some numbers have square roots which are airrational - that is they are not whole numbers nor can they be written as a fraction eg 6 - if you try to find the square root of 6 on your calculator you''ll find it will give you a decimal number to 8 decimal places but that decimal actually carries on infinitely

OpenStudy (cwrw238):

yes its not hard

OpenStudy (cwrw238):

these irrational roots are written as sqrt 6 , sqrt 19 etc